Celsius temperature readings can be converted to farenheit readings with the formual F=9/5C +32. What is the farenheit temperture that corresponds to each of the following celsius temperatures: -10,0,15,100?
Answer by Nate(3500)

F=(9/5)C +32
F=(9/5)(-10) +32
F=-18 +32
F=14
For -10 degrees celcius, the temperature is 14 degrees fahreheit.
F=(9/5)C +32
F=(9/5)0 +32
F=32
For 0 degrees celcius, the temperature is 32 degrees fahreheit.
F=(9/5)C +32
F=(9/5)(15) +32
F=27 +32
F=59
For 15 degrees celcius, the temperature is 59 degrees fahreheit.
F=(9/5)C +32
F=(9/5)(100) +32
F=180 +32
F=112
For 100 degrees celcius, the temperature is 212 degrees fahreheit.